OpenStudy (anonymous):

a truck covers 40 meters in a 8.50 seconds while smoothly slowing down to a final velocity of 2.80 meters per second a)find the trucks original speed b) find its acceleration

OpenStudy (shamim):

given final velocitu v=2.8m/s initial velocity u=? distance s=40 m time t=8.50 sec acceleration a=?

OpenStudy (anonymous):

You need to solve for two variables (V0 and A) so need at least two independent equations. There are two basic equations that you want to apply: (1) for acceleration: A = dV/dt = (V1 - V0)/8.5, with V1 is 2.8 m/s (2) for distance: Xt = X0 + V0*t + A/2*t^2 Fill out the second one first: 40 = 0 + V0*8.5 + A/2* 8.5^2, which still leaves you with two variables. Now use the first one, which relates one variable to the other. Replace A in the filled out second equation by the first equation. What you now get is a single equation with only one variable, V0, which is the initial velocity. Solve this and find V0. Use the first equation again to calculate A now you have found V0 already. See if you can calculate yourself....

OpenStudy (anonymous):

The numbers I found are -0.45 and 6.6 and I leave it up to you to determine which one is which.

OpenStudy (anonymous):

I keep on getting -0.22 I don't understand what I'm doubg wrong

OpenStudy (anonymous):

Doing*

OpenStudy (anonymous):

Hey Chicharito, can you describe your calculations here, so we can check. The order of your solution seems good, so you're on the right track.

OpenStudy (anonymous):